OpenStudy (anonymous):

how do I find the zeros of t^3 -3t algebraically

OpenStudy (acxbox22):

t^3-3t=0 ????<--- is that the equation

OpenStudy (anonymous):

yup

OpenStudy (acxbox22):

factor out the t t(t^2-3)=0 then t=0 and t^2-3=0 t^2=3 take the square root of both sides

OpenStudy (anonymous):

thanks

OpenStudy (acxbox22):

\(\Huge\color{red}{\ddot\smile}\Large\color{red}{NP}\)